OpenStudy (anonymous):

Can someone explain to me how to get my answer Evaluate and leave the solution in standard notation: (2 x 104)(5.1 x 102) A. 10.2 B. 10,200,000 C. 1,200,000 D. 102,000

OpenStudy (anonymous):

I am assuming that 104 and 102 mean 10^4 and 10^2 =2*10^4*5.1*10^2 =2*5.1*10^4+2 =10.2*10^6 =10200000 B.

OpenStudy (anonymous):

Yes sorry , so just multiply straight across ?

OpenStudy (anonymous):

Basically.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No prob.